OpenStudy (anonymous):

i need to find the derivative of f(x)=x^4(1-(2/x+1))

OpenStudy (amistre64):

the simplest way is to expand it all out to a sum of terms

OpenStudy (amistre64):

f(x)=x^4[1-(2/x+1)] f(x)= x^4 - 2x^4/(x+1) [f(x)= x^4 - 2x^4/(x+1)]' [f(x)]' = f'(x) [x^4]' = 4x^3 [- 2x^4/(x+1)]' = (x+1)[-2x^4]' - (-2x^4)[x+1]' ------------------------- (x+1)^2 -8x^4 -8x^3 +2x^4 = ------------------- (x+1)^2

OpenStudy (amistre64):

-6x^4 -8x^3 f'(x) = 4x^3 + ------------ (x+1)^2 4x^3(x+1)^2 -6x^4 -8x^3 f'(x) = ------------------------- (x+1)^2 4x^3(x^2+2x+1) -6x^4 -8x^3 f'(x) = --------------------------- (x+1)^2 4x^5 +8x^4 +4x^3 -6x^4 -8x^3 f'(x) = ----------------------------- (x+1)^2 \(4x^5 +2x^4 -4x^3\) f'(x) = --------------- \((x+1)^2\)

OpenStudy (amistre64):

if I kept track of it that shuld be good

OpenStudy (anonymous):

oh my gosh that is awesome help thank you so much!